OpenStudy (anonymous):

The y-intercept of h(x)=(x^2-16)/(x-4) is a. (0,4) b. (0,0) c. (0,-4) )

OpenStudy (anonymous):

Set x to zero and solve

OpenStudy (anonymous):

It doesnt give you x...

OpenStudy (anonymous):

\[\frac{x^2-16 }{x- 4 } = \frac{ (0)^2 -16 }{ 0-4} \]

OpenStudy (anonymous):

what it is asking when it says that y intercept, is when does y cross x.

OpenStudy (anonymous):

so my answer will be "c" because thats what I have **crosses fingers.

OpenStudy (anonymous):

nooo

OpenStudy (anonymous):

-16/-4 = 4 lol negs cancel

OpenStudy (anonymous):

(O,4) is your solution.

OpenStudy (anonymous):

Thank you... So basically i have to follow the rules of sign numbers with this?

OpenStudy (anonymous):

All you have to understand with this is that when it asks for the Y-intercept, what it wants to know is when the x value is zero, what is y. and then basic math skills

OpenStudy (anonymous):

Alright. Thank you.

OpenStudy (anonymous):

Yeah as in a graph, when the middle line of the t is zero. if that makes more sense.

OpenStudy (anonymous):